# Complexity transitions in chaotic quantum systems: Nonstabilizerness, entanglement, and fractal dimension in SYK and random matrix models

### v2

@@ I Introduction
The interplay between interactions and disorder Vojta (2019) in quantum many-body systems gives rise to complex systems with extremely rich phenomenology, characterized by varying degrees of chaoticity and ergodicity Gornyi and Mirlin (2002); Pino et al. (2016); Basko et al. (2006); Pino et al. (2016); Green (2015); Abanin et al. (2019). Although analytical approaches are available in certain limits of ergodic Berry (1977); Bohigas et al. (1984); Haake (1991); Lahoche and Ousmane Samary (2020) and localized regimes Evers and Mirlin (2008); Hauke and Heyl (2015); Altman and Vosk (2015); Goremykina et al. (2019), and while quantum simulations Hauke et al. (2012); Georgescu et al. (2014); Daley et al. (2022); Fraxanet et al. (2023) have explored various regimes of disordered interacting systems Jurcevic et al. (2014); Schreiber et al. (2015); yoon Choi et al. (2016); Smith et al. (2016); Bordia et al. (2017); Lüschen et al. (2017); Rispoli et al. (2019); Sauerwein et al. (2023), their treatment remains a significant challenge both for theory and experiment. In order to translate that difficulty into a measurable quantity, various notions have been proposed Osborne (2012); Orús (2019); Baiguera et al. (2025); Huang et al. (2024). Notable examples of such “complexity markers” include entanglement entropy Horodecki et al. (2009) and stabilizer Rényi entropy Leone et al. (2022), which have clear implications both for the difficulty of numerically treating Vidal (2003); Eisert (2013); Orús (2014, 2019); Montangero et al. (2018); Mora and Briegel (2005); Bravyi et al. (2019); Chernyshev et al. (2024); Brökemeier et al. (2025) the system under study as well as for the quantum resources Chitambar and Gour (2019) an experiment needs in order to generate the target state Joshi et al. (2023); Howard and Campbell (2017). Such markers are also deeply connected to the question of what actually defines a quantum system as complex Liu et al. (2018); Baiguera et al. (2024); Bu et al. (2024); Haug et al. (2024); Leone et al. (2023). As a disordered quantum many-body system enters into respectively less ergodic and more localized regimes, the number of basis states involved is reduced (see Fig. 1b), suggesting intuitively a suppression of quantum complexity. As of yet, it remains, however, an open question how precisely ergodicity transitions are related to changes in complexity markers, and in particular, whether these define qualitatively different complexity regimes.
@@
In this paper, we map out complexity quantifiers (see Fig. 1) in prototypical disordered models that exhibit multiple distinct regimes: the Rosenzweig–Porter (RP) model Rosenzweig and Porter (1960); Altland et al. (1997), the Power-Law Random Banded Matrix (PLRBM) model Mirlin et al. (1996), and the Sachdev–Ye–Kitaev (SYK) Sachdev and Ye (1993); Kitaev (2015) model coupled to an Ising chain (see Fig. 2). The RP and PLRBM are paradigmatic models of chaotic systems that interpolate between ergodic and non-ergodic regimes, and they are central to the understanding of Anderson localization Anderson (1958); Evers and Mirlin (2000) and (multi)fractal states Evers and Mirlin (2008); Kravtsov et al. (2015); Mirlin and Evers (2000). The SYK model is a prototypical example of a maximally chaotic quantum many-body system Maldacena et al. (2016a), which plays a key role in the phenomenological description of non-Fermi liquid behavior Sachdev and Ye (1993); Chowdhury et al. (2022) and in the study of quantum black holes through the holographic principle Sachdev (2015); Cotler et al. (2017). We scrutinize these models using three different markers of complexity: (i) The fractal dimension Castellani and Peliti (1986); Halsey et al. (1986); Aoki (1986); Kramer and MacKinnon (1993); Čadež et al. (2024), a key observable for identifying localized regions and mobility edges, quantifies how much a state is spread over a computational basis. (ii) The entanglement entropy across half–half bipartitions estimates a state’s quantumness in terms of its inability to be expressed as a direct product state Orús (2019). However, not all entangled states are computationally challenging. An outstanding example is the class of stabilizer states, which contain highly-entangled states but can be tackled efficiently classically thanks to the Gottesman–Knill theorem Gottesman (1998a, b). In contrast, non-stabilizer states fall outside this framework and are exponentially hard to represent on a classical computer. To capture this distinction, we use as the third measure (iii) the non-stabilizerness Veitch et al. (2014) (or “magic”) of a state, quantified through the stabilizer Rényi entropy (SRE) Leone et al. (2022). While the fractal dimension and entanglement entropy have been instrumental in characterizing non-ergodic quantum systems already for some time Mirlin et al. (1996); Refael and Moore (2009); Lukin et al. (2019); Sierant et al. (2025), magic is currently emerging as a powerful tool in various contexts, including the study of ergodicity Vairogs and Yan (2024); Chen et al. (2024a); Zhang and Gu (2024); Turkeshi et al. (2023, 2025a, 2025b); Falcão et al. (2025), quantum chaos Leone et al. (2021); Goto et al. (2022); Varikuti et al. (2025), and as a sensitive probe for identifying quantum phase transitions White et al. (2021); Oliviero et al. (2022); Tarabunga et al. (2023); Falcão et al. (2024); Jasser et al. (2025); Ding et al. (2025); Hoshino et al. (2025); Bera and Schirò (2025).
@@
In our work, we find that all three complexity markers show sharp phase transitions, separating regimes of different levels of complexity. Notably, the phase transition points are distinct when an extended fractal phase is present, as in the bulk of the RP spectrum. In contrast, when such a phase is absent, as in the ground state of the RP model, all considered markers detect the transition points close to each other, see Fig. 1c. Furthermore, an analysis of the SYK4+Ising model reveals that the SRE exhibits a richer structure than the other markers and, in particular, is significantly more sensitive to underlying symmetries of the many-body systems, such as fermion parity and time reversal. Overall, this comparative analysis suggests that no single marker can independently describe all relevant regimes of the system, which may undergo distinct complexity transitions. These results not only yield insights into the physics of disordered quantum many-body systems but also help identify the complexity marker best suited to capture the computational hardness across different phases Yang et al. (2017); Ghosh et al. (2023); Bu et al. (2024); Chernyshev et al. (2024); Brökemeier et al. (2025), guiding the selection of appropriate simulation methods.
@@
Finally, our studies complement ongoing research on several complexity markers in ergodic and integrable/localized systems, including Krylov complexity Rabinovici et al. (2022); Basu et al. (2024); Cohen et al. (2024); Sahu et al. (2024); Bhattacharjee and Nandy (2025), logarithmic multifractality Chen et al. (2024b), operator spreading Nahum et al. (2018), holographic complexity Alishahiha (2015), or quantum-computational complexity Watrous (2008).
@@
The remainder of the paper is structured as follows. In Sec. II, we present the complexity markers used to characterize the various quantum phases explored in this study. In Secs. III and IV, we apply these markers to analyze random matrix models, specifically the Rosenzweig–Porter model and the power-law random banded matrix model. In Sec. V, we investigate the Sachdev–Ye–Kitaev model interpolated towards the integrable Ising model. Our main findings and possible future research directions are presented in Sec. VI. In several Appendices, we present technical details on the calculations as well as supplementary findings, e.g., on a mobility edge in the RP model (Appendix A), on transitions in the self-averaging properties of the considered complexity markers (Appendix D), and on the Pauli spectrum of RP and PLRBM models (Appendix F).
@@ II.1 Fractal dimension
where $L=2^{N}$ is the Hilbert space dimension for a space of $N$ qubits. The fractal dimension $D_{q}$ is defined from the IPR as Halsey et al. (1986); Aoki (1986); Kramer and MacKinnon (1993)
@@
When the wavefunction $\psi$ is confined to a small region of the Hilbert space, the system is considered to be in the localized phase, marked by $D_{q}=0$ in the thermodynamic limit. For a fully extended state, in contrast, where the wavefunction spreads uniformly over all basis states, one obtains $D_{q}=1$. States in the ergodic phase, however, are usually not entirely uniformly distributed. Instead, they can be approximated by Haar-random states, which do not saturate the fractal dimension for finite system sizes. For $q=2$, the corresponding fractal dimension is given by Bäcker et al. (2019)
@@
For $N\to\infty$, one recovers $D_{2}^{\rm Haar}\to 1$. For non-extended states, as are expected to occur in non-ergodic but delocalized phases, just a fraction of the volume will be occupied, and we will have $D_{q}<1$. If $D_{q}$ is a constant across all values of $q$, the state is called single-fractal (or just fractal), which is the case for the Rosenzweig–Porter model in the non-ergodic extended regime Kravtsov et al. (2015). In the general case, $D_{q}$ is a function of $q$. Such a state is called multi-fractal. An example occurs in the power-law random banded model at the Anderson transition Mirlin et al. (1996); Evers and Mirlin (2000). In this paper, we will focus on $q=2$ only.
II.2 Entanglement entropy
The second marker we consider is the entanglement entropy maximized with respect to all possible equal-sized (half–half) bipartitions of the state in the computational basis. The maximal entanglement entropy is a key measure of quantum correlations between subsystems. The presence of entanglement has been identified as one of the first limitations for efficient classical simulations of quantum states Vidal (2003); Hastings (2007); Eisert et al. (2010), and entanglement entropy provides an estimate of the bond dimension required in tensor-network methods Eisert (2013); Orús (2014); Montangero et al. (2018); Orús (2019). Moreover, it is widely used to probe many-body localization and quantum phase transitions Wu et al. (2004); Osterloh et al. (2002); Amico et al. (2008); Calabrese and Cardy (2009); Geraedts et al. (2016).
@@ II.2 Entanglement entropy
For Haar-random states, the entanglement entropy follows the Page curve Page (1993), which for $L_{A}\leq L_{B}$ reads
@@ II.3 Stabilizer Rényi entropy (SRE)
In this work, we will use the second-order Rényi entropy, $q=2$. The SRE vanishes for a stabilizer state and grows as the state exhibits increasing degrees of magic. For Haar-random states, we have Turkeshi et al. (2025a) (see Appendix E for details),
@@
In what follows, we investigate how the three complexity markers discussed above evolve as a system transitions from a chaotic random matrix regime to a localized phase. We consider three representative models: the generalized Rosenzweig–Porter (RP) model (Sec. III), the Power-Law Random Banded Matrix (PLRBM) model (Sec. IV), and the Sachdev–Ye–Kitaev (SYK4) model coupled to an Ising chain (Sec. V). These models are characterized by a single tunable parameter, which controls the interpolation between chaotic and localized behavior. Moreover, the RP and PLRBM models exhibit Anderson criticality and support multifractal eigenstates Evers and Mirlin (2008); Kravtsov et al. (2015), providing a fertile ground for analyzing the distinct responses of the complexity markers.
III Rosenzweig–Porter model
@@
III.1 Model
The RP model is an ensemble of $L\times L$ random Hermitian matrices defined as Rosenzweig and Porter (1960)
@@ III.1 Model
The relative simplicity of the RP model enables an analytical treatment of both its spectral properties  Kunz and Shapiro (1998); Kravtsov et al. (2015) and eigenvector statistics  von Soosten and Warzel (2019); Bogomolny and Sieber (2018a); Truong and Ossipov (2016), allowing for a comprehensive characterization of the model in the bulk of the energy spectrum. In the ergodic phase, which spans the range $0\leq\gamma<1$, the model behaves like an RMT, with Wigner–Dyson level spacing and Porter–Thomas statistics for the eigenvector probability distribution, signaling quantum chaos and ergodicity in Hilbert space. At $\gamma=1$, a transition to non-ergodic behavior occurs, and at $\gamma=2$ Anderson localization sets in. In the regime $2<\gamma<\infty$, wavefunctions are localized within a small region of Hilbert space, and the level spacing distribution exhibits Poissonian statistics. A key feature of the RP model is the presence of an intermediate regime between the ergodic and localized phases, in the range $1<\gamma<2$, where wavefunctions remain extended but occupy only a fraction of the full Hilbert space, exhibiting fractal behavior that deviates from conventional ergodicity without being fully localized. The fractal dimension defined in Eq. (3) is a particularly effective marker for these transitions Kravtsov et al. (2015), leading to the characterization
@@
The same phase transitions have been confirmed not only through eigenvalue and eigenstate statistics but also via dynamical properties, such as diffusion characteristics Monthus (2017); Amini (2017), as well as survival and return probabilities Bera et al. (2018); Tomasi et al. (2019). Given the presence of different well-established regimes, the RP model serves as an ideal benchmark for evaluating how the behavior of the various complexity markers across different phases, as we discuss in detail in the following section.
@@
III.2 Results
In this section, we comparatively study the behavior of eigenvectors—both ground state and bulk states—undergoing a localization transition using the complexity markers defined in Sec. II: fractal dimension $\langle D_{2}\rangle$, maximum half-chain von-Neumann entanglement entropy $\langle S_{\rm vN}\rangle$, and stabilizer Rényi entropy $\langle\mathcal{M}_{2}\rangle$, where the angular brackets denote ensemble averaging. We first analyze the bulk eigenstates of the RP model in detail before extending our discussion to the ground state. We focus on the parameter range $\gamma\in[0,3]$ for varying system size $N=4$ to $12$ (corresponding to matrix dimension $2^{N}\times 2^{N}$), averaged over multiple samples.
@@ III.2 Results
Notably, the points $\gamma_{1}$ and $\gamma_{2}$, and therefore the parameter ranges corresponding to high- and low-complexity behavior, differ significantly between markers. For example, as shown in Fig. 3, the entanglement entropy features the broadest high-complexity plateau (panel b1), while the fractal dimension shows the narrowest (panel a1). We can make the separation into distinct regimes more precise by examining the first derivative of each marker with respect to $\gamma$. The intersection of these derivatives for different system sizes $N$ provides a reliable signature of emerging non-analytic behavior, indicating the onset of a sharp transition in the thermodynamic limit. We extract the transition points by extrapolating the crossing positions between two consecutive system sizes, $N_{1}$ and $N_{2}$, as a function of the inverse geometric mean, $1/\sqrt{N_{1}N_{2}}$ Pino et al. (2019). These extrapolations are shown in the insets of Fig. 3a2-c2.
@@
Among all the transition points, the most notable contrast lies in $\gamma_{1}^{\infty}$, where $S_{\rm vN}$ and $\mathcal{M}_{2}$ display a new phase transition that occurs significantly later than $\gamma=1$ and that is undetected by $D_{2}$. A similar trend is evident in their self-averaging behavior (see Appendix D), which undergoes abrupt changes that reinforce this observation. This behavior may stem from the fact that fractal states, while non-ergodic, can still exhibit high entanglement and high magic Turkeshi et al. (2023), causing these markers to remain large even after the system departs from full ergodicity. Consequently, the extended regime between $\gamma_{1}^{\infty}[D_{2}]$ and $\gamma_{1}^{\infty}[\mathcal{M}_{2}]$ represents a distinct non-ergodic phase, where the states remain highly entangled and highly magical, yet are no longer fully ergodic. This aligns with the intuition that the non-ergodic fractal phase can host highly complex states.
@@
The analysis of random matrix models is traditionally focused on the properties of the full eigenspectrum. However, their ground state tends to be more sensitive to ergodicity breaking, giving rise to a mobility edge  Basko et al. (2006); Sodin (2010); Chanda et al. (2020); Guo et al. (2021); Geißler and Pupillo (2020); Wei et al. (2020); Yousefjani and Bayat (2023) and motivating a dedicated analysis.
@@
IV Power-Law Random Banded Matrix model

### v3

@@ 1 Introduction
The interplay between interactions and disorder [1] in quantum many-body systems gives rise to complex systems with extremely rich phenomenology, characterized by varying degrees of chaoticity and ergodicity [2, 3, 4, 3, 5, 6]. Although analytical approaches are available in certain limits of ergodic [7, 8, 9, 10] and localized regimes [11, 12, 13, 14], and while quantum simulations [15, 16, 17, 18] have explored various regimes of disordered interacting systems [19, 20, 21, 22, 23, 24, 25, 26], their treatment remains a significant challenge both for theory and experiment. In order to translate that difficulty into a measurable quantity, various notions have been proposed [27, 28, 29, 30]. Notable examples of such “complexity markers” include entanglement entropy [31] and stabilizer Rényi entropy [32], which have clear implications both for the difficulty of numerically treating [33, 34, 35, 28, 36, 37, 38, 39, 40] the system under study as well as for the quantum resources [41] an experiment needs in order to generate the target state [42, 43]. Such markers are also deeply connected to the question of what actually defines a quantum system as complex [44, 45, 46, 47, 48]. As a disordered quantum many-body system enters into respectively less ergodic and more localized regimes, the number of basis states involved is reduced (see Fig. 1b), suggesting intuitively a suppression of quantum complexity. As of yet, it remains, however, an open question how precisely ergodicity transitions are related to changes in complexity markers, and in particular, whether these define qualitatively different complexity regimes.
@@
In this paper, we map out complexity quantifiers (see Fig. 1) in prototypical disordered models that exhibit multiple distinct regimes: the Rosenzweig–Porter (RP) model [49, 50], the Power-Law Random Banded Matrix (PLRBM) model [51], and the Sachdev–Ye–Kitaev (SYK) [52, 53] model coupled to an Ising chain (see Fig. 2). The RP and PLRBM are paradigmatic models of chaotic systems that interpolate between ergodic and non-ergodic regimes, and they are central to the understanding of Anderson localization [54, 55] and (multi)fractal states [11, 56, 57]. The SYK model is a prototypical example of a maximally chaotic quantum many-body system [58], which plays a key role in the phenomenological description of non-Fermi liquid behavior [52, 59] and in the study of quantum black holes through the holographic principle [60, 61]. We scrutinize these models using three different markers of complexity: (i) The fractal dimension [62, 63, 64, 65, 66], a key observable for identifying localized regions and mobility edges, quantifies how much a state is spread over a computational basis. (ii) The entanglement entropy across half–half bipartitions estimates a state’s quantumness in terms of its inability to be expressed as a direct product state [28]. However, not all entangled states are computationally challenging. An outstanding example is the class of stabilizer states, which contain highly-entangled states but can be tackled efficiently classically thanks to the Gottesman–Knill theorem [67, 68]. In contrast, non-stabilizer states fall outside this framework and are exponentially hard to represent on a classical computer. To capture this distinction, we use as the third measure (iii) the non-stabilizerness [69] (or “magic”) of a state, quantified through the stabilizer Rényi entropy (SRE) [32]. While the fractal dimension and entanglement entropy have been instrumental in characterizing non-ergodic quantum systems already for some time [51, 70, 71, 72], magic is currently emerging as a powerful tool in various contexts, including the study of ergodicity [73, 74, 75, 76, 77, 78, 79], quantum chaos [80, 81, 82], and as a sensitive probe for identifying quantum phase transitions [83, 84, 85, 86, 87, 88, 89, 90].
@@
In our work, we find that all three complexity markers show sharp phase transitions, separating regimes of different levels of complexity. Notably, the phase transition points are distinct when an extended fractal phase is present, as in the bulk of the RP spectrum. In contrast, when such a phase is absent, as in the ground state of the RP model, all considered markers detect the transition points close to each other, see Fig. 1c. Furthermore, an analysis of the SYK4+Ising model reveals that the SRE exhibits a richer structure than the other markers and, in particular, is significantly more sensitive to underlying symmetries of the many-body systems, such as fermion parity and time reversal. Overall, this comparative analysis suggests that no single marker can independently describe all relevant regimes of the system, which may undergo distinct complexity transitions. These results not only yield insights into the physics of disordered quantum many-body systems but also help identify the complexity marker best suited to capture the computational hardness across different phases [91, 92, 46, 39, 40], guiding the selection of appropriate simulation methods.
@@
Finally, our studies complement ongoing research on several complexity markers in ergodic and integrable/localized systems, including Krylov complexity [93, 94, 95, 96, 97], logarithmic multifractality [98], operator spreading [99], holographic complexity [100], or quantum-computational complexity [101].
@@
The remainder of the paper is structured as follows. In Sec. 2, we present the complexity markers used to characterize the various quantum phases explored in this study. In Secs. 3 and 4, we apply these markers to analyze random matrix models, specifically the Rosenzweig–Porter model and the power-law random banded matrix model. In Sec. 5, we investigate the Sachdev–Ye–Kitaev model interpolated towards the integrable Ising model. Our main findings and possible future research directions are presented in Sec. 6. In several Appendices, we present technical details on the calculations as well as supplementary findings, e.g., on a mobility edge in the RP model (Appendix A), on transitions in the self-averaging properties of the considered complexity markers (Appendix D), and on the Pauli spectrum of RP and PLRBM models (Appendix F).
@@ 2.1 Fractal dimension
where $L=2^{N}$ is the Hilbert space dimension for a space of $N$ qubits. The fractal dimension $D_{q}$ is defined from the IPR as [63, 64, 65]
@@
When the wavefunction $\psi$ is confined to a small region of the Hilbert space, the system is considered to be in the localized phase, marked by $D_{q}=0$ in the thermodynamic limit. For a fully extended state, in contrast, where the wavefunction spreads uniformly over all basis states, one obtains $D_{q}=1$. States in the ergodic phase, however, are usually not entirely uniformly distributed. Instead, they can be approximated by Haar-random states, which do not saturate the fractal dimension for finite system sizes. For $q=2$, the corresponding fractal dimension is given by [104]
@@
For $N\to\infty$, one recovers $D_{2}^{\rm Haar}\to 1$. For non-extended states, as are expected to occur in non-ergodic but delocalized phases, just a fraction of the volume will be occupied, and we will have $D_{q}<1$. If $D_{q}$ is a constant across all values of $q$, the state is called single-fractal (or just fractal), which is the case for the Rosenzweig–Porter model in the non-ergodic extended regime [56]. In the general case, $D_{q}$ is a function of $q$. Such a state is called multi-fractal. An example occurs in the power-law random banded model at the Anderson transition [51, 55]. In this paper, we will focus on $q=2$ only.
2.2 Entanglement entropy
The second marker we consider is the entanglement entropy maximized with respect to all possible equal-sized (half–half) bipartitions of the state in the computational basis. The maximal entanglement entropy is a key measure of quantum correlations between subsystems. The presence of entanglement has been identified as one of the first limitations for efficient classical simulations of quantum states [33, 105, 106], and entanglement entropy provides an estimate of the bond dimension required in tensor-network methods [34, 35, 36, 28]. Moreover, it is widely used to probe many-body localization and quantum phase transitions [107, 108, 109, 110, 111].
@@ 2.2 Entanglement entropy
For Haar-random states, the entanglement entropy follows the Page curve [112], which for $L_{A}\leq L_{B}$ reads
@@ 2.3 Stabilizer Rényi entropy (SRE)
In this work, we will use the second-order Rényi entropy, $q=2$. The SRE vanishes for a stabilizer state and grows as the state exhibits increasing degrees of magic. For Haar-random states, we have [77] (see Appendix E for details),
@@
In what follows, we investigate how the three complexity markers discussed above evolve as a system transitions from a chaotic random matrix regime to a localized phase. We consider three representative models: the generalized Rosenzweig–Porter (RP) model (Sec. 3), the Power-Law Random Banded Matrix (PLRBM) model (Sec. 4), and the Sachdev–Ye–Kitaev (SYK4) model coupled to an Ising chain (Sec. 5). These models are characterized by a single tunable parameter, which controls the interpolation between chaotic and localized behavior. Moreover, the RP and PLRBM models exhibit Anderson criticality and support multifractal eigenstates [11, 56], providing a fertile ground for analyzing the distinct responses of the complexity markers.
3 Rosenzweig–Porter model
@@
3.1 Model
The RP model is an ensemble of $L\times L$ random Hermitian matrices defined as [49]
@@ 3.1 Model
The relative simplicity of the RP model enables an analytical treatment of both its spectral properties  [120, 56] and eigenvector statistics  [121, 122, 123], allowing for a comprehensive characterization of the model in the bulk of the energy spectrum. In the ergodic phase, which spans the range $0\leq\gamma<1$, the model behaves like an RMT, with Wigner–Dyson level spacing and Porter–Thomas statistics for the eigenvector probability distribution, signaling quantum chaos and ergodicity in Hilbert space. At $\gamma=1$, a transition to non-ergodic behavior occurs, and at $\gamma=2$ Anderson localization sets in. In the regime $2<\gamma<\infty$, wavefunctions are localized within a small region of Hilbert space, and the level spacing distribution exhibits Poissonian statistics. A key feature of the RP model is the presence of an intermediate regime between the ergodic and localized phases, in the range $1<\gamma<2$, where wavefunctions remain extended but occupy only a fraction of the full Hilbert space, exhibiting fractal behavior that deviates from conventional ergodicity without being fully localized. The fractal dimension defined in Eq. (3) is a particularly effective marker for these transitions [56], leading to the characterization
@@
The same phase transitions have been confirmed not only through eigenvalue and eigenstate statistics but also via dynamical properties, such as diffusion characteristics [124, 125], as well as survival and return probabilities [126, 127]. Given the presence of different well-established regimes, the RP model serves as an ideal benchmark for evaluating how the behavior of the various complexity markers across different phases, as we discuss in detail in the following section.
@@
3.2 Results
In this section, we comparatively study the behavior of eigenvectors—both ground state and bulk states—undergoing a localization transition using the complexity markers defined in Sec. 2: fractal dimension $\langle D_{2}\rangle$, maximum half-chain von-Neumann entanglement entropy $\langle S_{\rm vN}\rangle$, and stabilizer Rényi entropy $\langle\mathcal{M}_{2}\rangle$, where the angular brackets denote ensemble averaging. We first analyze the bulk eigenstates of the RP model in detail before extending our discussion to the ground state. We focus on the parameter range $\gamma\in[0,3]$ for varying system size $N=4$ to $12$ (corresponding to matrix dimension $2^{N}\times 2^{N}$), averaged over multiple samples.
@@ 3.2 Results
Notably, the points $\gamma_{1}$ and $\gamma_{2}$, and therefore the parameter ranges corresponding to high- and low-complexity behavior, differ significantly between markers. For example, as shown in Fig. 3, the entanglement entropy features the broadest high-complexity plateau (panel b1), while the fractal dimension shows the narrowest (panel a1). We can make the separation into distinct regimes more precise by examining the first derivative of each marker with respect to $\gamma$. The intersection of these derivatives for different system sizes $N$ provides a reliable signature of emerging non-analytic behavior, indicating the onset of a sharp transition in the thermodynamic limit. We extract the transition points by extrapolating the crossing positions between two consecutive system sizes, $N_{1}$ and $N_{2}$, as a function of the inverse geometric mean, $1/\sqrt{N_{1}N_{2}}$ [128]. These extrapolations are shown in the insets of Fig. 3a2-c2.
@@
Among all the transition points, the most notable contrast lies in $\gamma_{1}^{\infty}$, where $S_{\rm vN}$ and $\mathcal{M}_{2}$ display a new phase transition that occurs significantly later than $\gamma=1$ and that is undetected by $D_{2}$. A similar trend is evident in their self-averaging behavior (see Appendix D), which undergoes abrupt changes that reinforce this observation. This behavior may stem from the fact that fractal states, while non-ergodic, can still exhibit high entanglement and high magic [76], causing these markers to remain large even after the system departs from full ergodicity. Consequently, the extended regime between $\gamma_{1}^{\infty}[D_{2}]$ and $\gamma_{1}^{\infty}[\mathcal{M}_{2}]$ represents a distinct non-ergodic phase, where the states remain highly entangled and highly magical, yet are no longer fully ergodic. This aligns with the intuition that the non-ergodic fractal phase can host highly complex states.
@@
The analysis of random matrix models is traditionally focused on the properties of the full eigenspectrum. However, their ground state tends to be more sensitive to ergodicity breaking, giving rise to a mobility edge  [4, 129, 130, 131, 132, 133, 134] and motivating a dedicated analysis.
@@
4 Power-Law Random Banded Matrix model